OpenStudy (emilyjones284):

help please... ill give a medal (:

OpenStudy (anonymous):

What's question?

OpenStudy (emilyjones284):

\[2^\frac{ 1 }{ 3}\times4^\frac{ 1 }{ 3 }\]

OpenStudy (emilyjones284):

simplify using the radical form

OpenStudy (emilyjones284):

the fractions are exponents by the way

OpenStudy (rsadhvika):

\(\large 2^\frac{1}{3} \times 4^\frac{1}{3}\)

OpenStudy (noelgreco):

Try this: \[2^{\frac{ 1 }{ 3 }}\times(2^{2})^{\frac{ 1 }{ 3 }}\] Once you have the same base use the rules of exponents.

OpenStudy (rsadhvika):

\(\large 2^\frac{1}{3} \times 4^\frac{1}{3}\) \(\large (2 \times 4)^{\frac{1}{3}} \) \(\large (2 \times 2^2)^\frac{1}{3}\) \(\large (2^3)^\frac{1}{3}\)

OpenStudy (emilyjones284):

thanks!

OpenStudy (rsadhvika):

yw :)